Find two consective integers such that the square of the sum of the two integers is 11 more than the smaller integer.
Answer by edjones(8007)

Let the integers be x & x+1
(x+x+1)^2=x+11
(2x+1)^2=x+11
4x^2+4x+1=x+11
4x^2+3x-10=0
4*-10=-40 what are the 2 factors of -40 whose sum = 3? 8 and -5
4x^2+8x-5x-10=0
4x(x+2)-5(x+2)=0
(4x-5)(x+2)=0
.
4x=5
x=5/4 not an integer
.
x=-2
x+1=-1
.
Ed